40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 9 11:46:01 2021, Anchal, Paco, Summary, AUX, Xend Green Laser PDH OLTF measurement image-6f2923a3-01ce-4d04-bc53-d8db0238e195.jpgimage-72223f4b-3b74-4574-a7ad-de6628a2c5e9.jpgX_Green_ARM_PDH_OLTF.pdf
    Reply  Thu Jun 10 14:01:36 2021, Anchal, Summary, AUX, Xend Green Laser PDH OLTF measurement loop algebra AUX_PDH_LOOP.pdf
       Reply  Mon Jun 14 18:57:49 2021, Anchal, Update, AUX, Xend is unbearably hot. Green laser is loosing lock in 10's of seconds XAUX_PDH_Err_In_ASD.pdfXAUX_PZT_Out_Mon_ASD.pdf
       Reply  Tue Jun 15 15:26:43 2021, Anchal, Paco, Summary, AUX, Xend Green Laser PDH OLTF measurement loop algebra, excitation at control point AuxPDHloop.pdf
          Reply  Fri Jun 18 10:07:23 2021, Anchal, Paco, Summary, AUX, Xend Green Laser PDH OLTF with coherence XEND_PDH_OLTF_with_Coherence.pdfBeta_Amp.pdf
Message ID: 16194     Entry time: Wed Jun 9 11:46:01 2021     Reply to this: 16197
Author: Anchal, Paco 
Type: Summary 
Category: AUX 
Subject: Xend Green Laser PDH OLTF measurement 

We measured the Xend green laser PDH Open loop transfer function by following method:

  • We first measured the feedback transfer function 'K' directly.
    • See attachment 2 for this measurement. We measured Out2/exc here.
  • Then, we closed the loop as shown in attachment 1with SR560 as a summing juntion at error point.
    • We injected excitation through B channel in SR560 and measured transfer function Out1/Out2.
    • This measurement should give us G_{OL} / K by loop alegbra.
  • Then we multiplied the two transfer function measurements to get open loop transfer function.

Result:

  • Our measurement gives the same UGF of 10kHz and phase margin of 53.5 degrees as reported in 13238.
  • The shape of measurement also follows 1/f above 10 Hz atleast.
  • Our measurement might not be correct below 10 Hz but we did not see any saturation or loss of lock in 1Hz to 10 Hz measurement.
  • This OLTF is different from the modelled OLTF here even though the UGF matches.
  • The feedback gain is supposed to roll-off faster than 1/f in 30Hz to 1kHz region but it does not seem to in our measurement.
  • This suggests that the actual uPDH box is shaping the loop different from what schematic suggests. This might mean that the gain is much lower in the low frequency region than we would like it to be.
  • We will investigate the reason of difference between model and measurement unless someone has a better explaination for the descripancy.
Attachment 1: image-6f2923a3-01ce-4d04-bc53-d8db0238e195.jpg  783 kB  Uploaded Wed Jun 9 13:08:04 2021  | Hide | Hide all
image-6f2923a3-01ce-4d04-bc53-d8db0238e195.jpg
Attachment 2: image-72223f4b-3b74-4574-a7ad-de6628a2c5e9.jpg  625 kB  Uploaded Wed Jun 9 13:08:56 2021  | Hide | Hide all
image-72223f4b-3b74-4574-a7ad-de6628a2c5e9.jpg
Attachment 3: X_Green_ARM_PDH_OLTF.pdf  30 kB  Uploaded Wed Jun 9 13:10:13 2021  | Hide | Hide all
X_Green_ARM_PDH_OLTF.pdf
ELOG V3.1.3-